OpenStudy (anonymous):

Walnuts cost $10 per pound and peanuts cost $5 per pound. How many pounds of peanuts would be needed for a 40 pound mixture that sells for $7.75 per pound? How would I set this up in a system of equations again??

OpenStudy (anonymous):

This is what I think it is- 10x+5y=7.75 x+y=7.75

OpenStudy (anonymous):

x + y = 40 I think..

OpenStudy (anonymous):

Okay I wasn't so sure if it was that or x+y=40. Thx I"ll try it out to see if I get an answer from my choices :)

OpenStudy (anonymous):

Their mixture is 40: So: x + y = 40 10x + 5y = 7.75 It is better to divide second equation by 5 first: 2x + y = 1.55 Now you have to just subtract them..

OpenStudy (anonymous):

Tyvm :) I got y=18 and x=22

OpenStudy (anonymous):

y=peanuts ; x=walnuts

OpenStudy (anonymous):

Welcome dear..